John's Furniture store sells customers an entire houseful

of furniture in one transaction. John has notice that his company

could use a bigger facility. There are two available locations.

Location A would have annual fixed costs of $120,000 and variable costs of $15,000 per furniture set; location B would have annual fixed costs of $320,000 and variable costs of $10,000 per furniture set. The furniture sets sell for $20,000.


a. For what output level would John be indifferent between the two locations?


  Output level    units


b-1.

For what output range would A be preferred? (Enter your answer as a whole number. Do not include the indifference point in your answer.)


Output range 0 to   


b-2.

For what output range would B be preferred? (Enter your answer as a whole number. Do not include the indifference point in your answer.)


  Output range or more
Answer #1

a)

for indifference total cost of location A =total cost of location B

120,000+(15000x=320000+10000x

5000x =200000

Output level =x=200000/5000=40 units

b-1)

Output range 0 to 40

b-2)

  Output range 40 or more